The probability that a first year student entering a certain private college needs neither a developmental math course nor a developmental English is 59% while 29% require a developmental math course and 31% require a developmental English course.
Find the probability that a first year student requires both a development math course and a developmental English course.

Answer by Boreal(15235)

P(M or E)=P(M)+P(E)-P(Both); probability one or the other is 1-P(neither)=0.41
0.41=0.29+0.31-P(both)
P(both required)=0.19, the answer